Two planes leave the airport at the same time. One flies directly east at 540 miles per hour. The other flies west at 720 miles per hour. How long will it take them to be 1890 miles apart?
Answer by Maths68(1474)

Effective Speed = 540+720 = 1260 mph
Distance = 1890 miles
Time = Distance / Speed
Time = 1890/1260
time = 1.5 hour.

After 1.5 hours two planes will be 1890 miles apart?


Check
======
Plane A Data
============
Speed = 540 mph
Time = 1.5 hour
Distance covered = ?
Distance = Speed * Time
Distance covered= 540 * 1.5 = 810 miles

Plane B Data
============
Speed = 720 mph
Time = 1.5 hour
Distance covered = ?
Distance = Speed * Time
Distance covered = 720 * 1.5 = 1080 miles

Total Distance between two planes after 1.5 hours = 1080+810 = 1890 miles